### v1

# SoK: Root Cause of $1 Billion Loss in Smart Contract Real-World Attacks via a Systematic Literature Review of Vulnerabilities

Abstract
The Ethereum ecosystem, despite its maturity, continues to witness catastrophic attacks, with billions of dollars in assets lost annually. In response, a significant body of research has focused on identifying and mitigating smart contract vulnerabilities. However, these efforts predominantly focus on implementation-level bugs, leaving a critical gap between academic understanding of vulnerabilities and the root causes of real-world high-impact financial losses. We employ a two-pronged methodology: first, a systematic literature review of 71 academic papers to build a comprehensive and up-to-date catalog of 24 active and 5 deprecated vulnerabilities as understood by the research community. Second, we conduct an in-depth, empirical analysis of 50 of the most severe real-world exploits between 2022 and 2025, collectively incurring over $1.09B in losses, to identify their true root causes. We introduce the concept of “exploit chains” by revealing that many incidents are not caused by isolated vulnerabilities but by combinations of human, operational, and economic design flaws that link with implementation bugs to enable an attack. Our analysis yields insights on how DApps are exploited in practice, leading to a novel, four-tier root-cause framework that moves beyond code-level vulnerabilities. We find that real-world successful attacks on Ethereum (and related networks) trace back to one of the four tiers of (1) protocol logic design, (2) lifecycle and governance, (3) external dependencies, and (4) traditional implementation bugs (classic smart contract vulnerabilities). We investigate the suitability of this multi-tier incident root-cause framework via a case study.
I Introduction
The Ethereum blockchain and its smart contract ecosystem have grown to manage hundreds of billions of dollars [1, 2]. Along with this growth, in recent years, a relentless series of successful high-profile hacks has occurred, resulting in staggering financial losses [3, 4]. Some reports suggest up to $ billion stolen from protocols in 2024 [5]. In response to such attacks, a substantial body of research has been performed to classify and detect smart contract vulnerabilities. These efforts focused on implementation-level bugs such as reentrancy and integer overflows [6, 7]. This led to the development of numerous static and dynamic analysis tools like Slither [8] and Echidna [9] or even execution frameworks to run a swarm of analysis tools on a contract, such as SmartBugs [10]. Despite their usefulness in detecting certain classes of vulnerabilities [11], both the severity (total funds stolen) and number of successful attacks have increased in recent years [5].
| Paper | Year | Vulns | Prev.b | Tax.a | Wildc | Key Strength |
| S1 [12] | 2020 | 40 | P | 4L | ✗ | Vulnerability attack mapping |
| S2 [13] | 2021 | 20 | ✗ | 3C | ✗ | Comprehensive tool analysis |
| S3 [14] | 2021 | 8 | ✗ | NI | C | Maps vulnerabilities attacks detection tools |
| S4 [15] | 2021 | 202 | ✓ | FM | ✗ | Comprehensive formal methods for all SCi verification. |
| S5 [16] | 2022 | 6 | ✗ | DT | ✗ | Vulnerability detection tools quality gaps |
| S6 [17] | 2022 | 24 | ✗ | RS | ✗ | Causal hierarchy enabling systematic prevention |
| S7 [18] | 2023 | 12 | ✗ | 3L | ✗ | Layer-based classification: Solidity, EVM, Blockchain |
| S8 [19] | 2023 | 6 | ✗ | TB | ✗ | Categorizes 6 vulns (BE/RC/EX/ED/BD)g + tool assessment |
| S9∗ [3] | 2023 | 50+ | ✓ | 5L | 181 | Analyzes 181 exploits across NET/CON/SC/PRO/AUXe layers |
| S10 [20] | 2024 | 94 | ✗ | HR | E | Open hierarchical taxonomy with code examples & ODCh |
| S11∗ [21] | 2024 | Vd | ✗ | UN | ✗ | ERj data model for unifying SCi vulnerability taxonomies |
| S12∗ [22] | 2024 | 45 | ✓ | 4L | 92 | Traces vulnerabilities in real attacks |
| S13 [23] | 2025 | 20+ | ✓ | 5L | ✗ | Comprehensive 5-layer attack classification |
| Our Work | 2025 | 24+5f | ✓ | Cause | 50 | Real-world validation, root-cause framework from incidents |
Footnotes: ∗SoK aTax: xL=x-Layer, xC=x-Category FM=Formal Methods, DT=Detection-based, RS=Root-cause-based, TB=Tool-Based, HR=Hierarchical, UN=Unified • bPrev: ✓=comprehensive, P=partial, ✗=none • cWild: Numbers=exploit count, C=case studies, E=examples, ✗=none • Vd=Various/unspecified count • eNET/CON/SC/PRO/AUX=Network/Consensus/Smart Contract/Protocol/Auxiliary • f core + 5 deprecated vulnerabilities • gBE=Bugs/Errors, RC=Rogue Contracts, EX=Extortion, ED=External Data, BD=Backdoors • hODC=Orthogonal Defect Classification • iSC=Smart Contract • jER=Entity-Relationship
Related SoKs and Surveys
Table I presents a review of three Systematization of Knowledge (SoK) and ten Systematic Literature Review (SLR) papers on Decentralized Application (DApp) security. Surveys and SoKs classify tens of vulnerabilities across different layers of the blockchain stack (e.g., network, smart contract, and protocol layers) [12, 17, 18]. These taxonomies have become increasingly granular in classifying weaknesses [17] and mapping them to formal verification properties [15], or aligning them with frameworks such as NIST bugs [14]. However, our review unveils a few gaps in this literature: First, prevention gap: only 4 of 13 works (S4, S9, S12, and S13) address prevention methods, most of which focus exclusively on detection. Second, the validation gap: merely 3 works validate their findings against real-world exploits—S3, S9 and S12 or analyze case studies, and exploits, while others rely on synthetic datasets. Third, the taxonomy evolution shows a shift from simple categorization to sophisticated multi-layer frameworks (4L, 5L), yet none adopt a root-cause approach based on real incidents.
Real-World Incidents Tell-Tale
At 08:56 UTC on 13 March 2023, the Euler Finance lending market was solvent; four minutes later, it was $ million in deficit [24, 25]. Nothing in its Solidity source code could be attributed to common smart contract vulnerabilities, no unchecked casts, no arithmetic overflow, or no reentrancy bug. Instead, an unguarded auxiliary function, a liquidation discount, and a flash loan combined into an unexpected profit engine to exploit the contract. Several months earlier, the Nomad bridge was emptied of $M because a single initializer function had been called with the wrong argument during an upgrade, resetting its trusted root to zero address and turning every message into valid [26]. Episodes like these are not rare accidents [5]. The vast majority of existing literature and automated tooling remains heavily focused on implementation-level weaknesses, the “known unknowns” [12, 17, 15]. Yet, as incidents like the $M Euler Finance exploit [24] show, financially devastating attacks may not merely stem from coding mistakes. In the mentioned protocol’s case, the attack succeeded as the liquidation mechanism was flawed. The code executed exactly as intended; it was the intention itself that created the vulnerability. This incident is not an isolated case. It highlights a systemic issue: while literature focuses on finding bugs in the code, a large-scale analysis of real-world attacks and comparing the exploited vulnerabilities in these incidents and their alignment with the literature is required.
1.
RQ1: Which vulnerabilities still affect Ethereum smart contracts as of , and how can they be mitigated?
2.
RQ2: What factors, other than mere coding vulnerability patterns (identified in RQ1), are evident in the incidents and contribute to the success of the attack?
3.
RQ3: What categories do these vulnerabilities fall into based on their root causes?
To address RQ1, we conduct a systematic literature review (SLR) of seminal papers, synthesizing their findings into a comprehensive catalog of active and deprecated vulnerabilities. This provides a robust, state-of-the-art baseline of academically recognized threats. To address RQ2, we perform a deep, qualitative analysis of 50 high-impact smart contract exploits that occurred between 2022 and early 2025, representing a total of $ in losses. By examining post-mortem reports, on-chain transaction data, and the attacked protocols’ source code and documentations, we move beyond bug labels (e.g., reentrancy) to uncover the steps (exploit chains Section IV) that enabled these attacks. This approach yielded a novel, four-tier root-cause framework derived directly from our empirical analysis, complementing the academic literature (Section V). We argue that real-world exploits are best understood as originating from one these four tiers: 1) flawed economic design and protocol logic (%), 2) protocol lifecycle and Governance failures (%), 3) external dependency vulnerabilities (%), and 4) implementation-level weaknesses (%).
In summary, this paper makes the following contributions:
1.
An up-to-date, curated catalog of active and deprecated smart contract vulnerabilities based on a systematic review of papers. For each active vulnerability type we include detailed prevention/mitigation strategies and real-world incidents involving (in the accompanying repository111https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents).
2.
A manual analysis of high-profile successful incidents222Distinct incidents from the our real-world examples in contribution 1. (selected from a pool of incidents) occurred in 2022–2025 on various protocol types (e.g., DeFi, cross-chain bridges, etc.) demonstrating that the most severe real-world exploits stem from factors beyond mere/isolated implementation mistakes. We identify exploit chains that combine human, operational, and economic design flaws with implementation bugs to enable attack vectors.
3.
A novel, four-tier root-cause framework for smart contract exploits, grounded on our analysis of real-world exploit methods.
4.
A case study on a sub-category of our root-cause-based framework with defense strategies against exploits targeting this sub-category.
5.
A dataset of real-world successful attack attributes, including exploited vulnerability types (confirmed by manual expert analysis), victim and target account addresses, and attack transaction hashes.
By systematizing the knowledge of how smart contracts fail in practice (and attributing these failures to vulnerabilities found in our SLR and root-causes in our four-tier framework), we show that many successful attacks stem not just from code-level bugs, but from a combination of design, governance, dependency, and operational factors. The rest of the paper is organized as follows. Section II presents our rigorous SLR and incident review protocol. Section III presents the results of our SLR for RQ1. Section IV presents the results of our real-world incident analysis insights for RQ2. Section V demonstrates our multi-tier incident root-cause framework for RQ3. Finally, Section VI concludes the paper.
II Methodology
To answer RQ1–RQ3, we analyzed both academic literature and high-quality technical smart contract incident (successful attack) reports. Sections II-A and II-B elaborate on the protocol used to identify, collect, and review these resources.
II-A Academic Literature Study Protocol
To answer RQ1, we conducted a systematic literature review (SLR) to obtain a comprehensive and structured understanding of Ethereum smart contract vulnerabilities identified in the academic literature. Following Kitchenham’s guidelines [27], our review process includes study identification (Section II-A1) and study selection (Section II-A2).
II-A1 Study Identification
Here, we first present the process of formulating search queries for scientific databases. Then, we present the process of filtering them to collect relevant scientific literature from reputable sources (briefly in Figure 1).
Search Keywords
We used the PICOC (Population, Intervention, Comparison, Outcome, Context) framework [27] to achieve a systematic search. According to this method, we defined keywords for our search in Table II. The “Population” was defined as the technology (smart contracts, blockchain), and the “Intervention” as the problem being investigated.
| Cat. | Keyword | Synonyms |
| P | “Ethereum smart contracts” | “Ethereum contracts”, “Solidity” |
| I | “Vulnerabilities” | “Security Flaws” OR “Security Weaknesses” |
| O | “Mitigation” | “Classification”, “Mitigation”, “Prevention”, “Categorization”, “classification”, “taxonomy” |
For search keywords, we added synonyms to the main keywords for varied searches. The final query for searching in scientific databases is the result of concatenating all keywords for P, I, and O (and their synonyms) from Table II.
Study Sources
We queried five databases: ACM, IEEE Xplore, Scopus, Springer, and Google Scholar (for pre-prints, theses, and gray literature) to collect studies. Figure 1 presents the number of papers retrieved from each: ACM (), IEEE (), Scopus (), Springer (), and Google Scholar ().
II-A2 Study Selection
Following the initial identification of papers, we removed duplicates and filtered them based on the exclusion criteria in Table III. Specifically, language (EC2) and publication year 2018–2024 (EC1) filtering reduced the count to papers. Title and abstract screening using criteria EC3 (Ethereum focus), EC5, and EC6 (topic relevance) yielded papers. Full-text assessment based on availability (EC4) further refined the selection to papers. Finally, quality-based evaluation via manual expert empirical validation and identification of novel contributions resulted in papers for in-depth analysis. Figure 1 illustrates this process.
II-B Real-World Incident Review Protocol
To answer RQ2, we reviewed post-mortem incident reports. To ensure the relevance of the successful hacks and maintain quality control over the reviewed resources, we applied a four-stage incident review protocol in Figure 2.
Stage 1: Comprehensive Data Collection. We aggregated incident reports from four reliable blockchain security sources: Rekt News [28], REKT database [29], SlowMist Hacked database [30], and BlockSec reports [31]. From each source, we extracted up to 100 of the most recent attacks based on the occurrence date. The report sources are selected for their report quality and analytical depth. In BlockSec’s case, our temporal parameters (April 2022 to April 2025) yielded only 37 documented attacks meeting our criteria, all of which were incorporated into our analysis. As a result, our initial dataset comprised a total of attacks for analysis.
Stage 2: Impact-Based Prioritization. Incidents underwent prioritization based on financial impact severity, emphasizing attacks resulting in substantial monetary losses. This criterion ensured our analysis concentrated on vulnerabilities manifesting the most significant ecosystem consequences.
Stage 3: Ecosystem Classification. We focused on incidents occurring within the Ethereum blockchain and its compatible networks. Each case was evaluated for relevance, resulting in a final dataset of distinct incidents that offer sufficient breadth while maintaining ecosystem coherence. We then mapped each incident to its respective blockchain environment, as shown in the network column of Table V, establishing three network categories:
1.
Ethereum Mainnet: The Ethereum network where native consensus exist, identified by transactions on etherscan.io ( attacks).
2.
Ethereum Layer 2 Solutions: These include Arbitrum, Optimism, Base, zkSync, Linea, Scroll, StarkNet, and Mode ( attacks). L2s inherit security from Ethereum while processing transactions off-chain [32]. They use either Optimistic or Zero-Knowledge rollups to post transaction data back to Layer 1.
3.
Ethereum Side-Chains: Networks like Polygon, BSC (Binance Smart Chain), and Ronin that maintain independent consensus mechanisms while supporting Ethereum Virtual Machine (EVM) compatibility [33] ( attacks).
This classification follows the technical distinction that true L2 solutions derive their security from Ethereum itself, while side-chains operate with separate validator sets and consensus mechanisms. For example, Base is considered an L2 because it uses the OP Stack to inherit Ethereum’s security properties, whereas Polygon PoS uses its own validator set for consensus, making it a side-chain.
| ID | Exclusion Criteria |
| EC1 | The article was published before 2018.a |
| EC2 | The article is written in a language other than English. |
| EC3 | The article focuses on vulnerabilities in platforms other than Ethereum. |
| EC4 | There is no full-text version of the article available. |
| EC5 | The article is a book chapter or editorial notes. |
| EC6 | The article does not concern any of these topics: survey of smart contract vulnerabilities, security of smart contracts, vulnerability detection or mitigation methods for smart contracts, demonstrates findings for smart contracts through simulation or empirical evidence. |
a
The vulnerabilities before are included via more recent papers.
Stage 4: Vulnerability Verification and Report Quality Assessment. In the final step, we applied two strict checks: (1) we made sure the incident involved at least one smart contract vulnerability not merely like infrastructure failure, human error, or stolen private keys; and (2) we checked that the incident was well-documented. To be included, a report had to clearly explain what the vulnerability was (including vulnerability localization information), how it was exploited, and what caused it. We excluded any cases that lacked enough technical detail. This process yielded a final dataset of well-documented, high-impact smart contract incidents with at least one smart contract vulnerability exploited in each incident. We use this dataset as the basis for our analysis (Section IV), which later resulted in a root cause-based four-tier classification of RQ3 (Section V). We cite a total number of reviewed reports for RQ2 and RQ3.
III Current Ethereum Smart Contract Vulnerabilities and Their Prevention/Mitigation
To answer RQ1, we conducted a comprehensive review of the literature (Section II-A), identifying distinct smart contract vulnerabilities. Upon further analysis, we determined that of these vulnerabilities have become obsolete by protocol-level upgrades (Section III-B4 and Section III-B5), compiler improvements (Section III-B2 and Section III-B3), or re-contextualized based on widespread usage (see Section III-B1). This left active smart contract vulnerabilities based on the academic literature. For each of the active vulnerabilities, we collected prevention and mitigation strategies from a combination of academic sources and technical reports confirmed by expert review (Section III-A).
III-A Active Vulnerabilities
Our SLR resulted in the conclusion that a core set of vulnerabilities remains exploitable and economically attractive. We unpack each of these active threats here. For every vulnerability we (i) recap a real-world incident or a proof-of-concept vulnerability (full information is included in our companion repository), (ii) expose the coding pattern that enables it, and (iii) distill the prevention or mitigation techniques. Taken together, this section provides a current, practitioner-oriented catalog for anyone designing, reviewing, or formally verifying Ethereum smart contracts.
III-A1 Integer Over/Underflow and Rounding Errors
These vulnerabilities stem from Solidity’s limitations in handling of numerical data types. Integer Over/Underflow vulnerability arises when the result of a calculation exceeds the lower or upper limits of the variable type size, making it impossible to be expressed within that type. This is common with smaller data types. For instance, in LABEL:lst:overflow, the balance stored in balances mapping will reset to zero if it reaches the unit value of the upper limit . Additionally, rounding errors occur when division operations are performed on integers in Solidity, which leads to inconsistent results of calculations, especially when precision is required [34].
In the listing LABEL:lst:overflow, the function forceOverflow attempts to assign the maximum possible value to a user’s balance and then increase it by one, which in earlier Solidity versions (<0.8.0) would have resulted in an integer overflow. Similarly, the withdraw function reduces the user’s balance, which in older versions, if unchecked, could lead to an integer underflow. Solidity 0.8.0 and later versions mitigate these issues by introducing built-in overflow and underflow checks, making such operations revert automatically.
Mitigation/Prevention
To prevent arithmetic vulnerabilities in Solidity smart contracts, developers commonly rely on SafeMath [35], built-in checks in Solidity , and compiler warnings. When an overflow or underflow occurs, the SafeMath library reverts the transaction. However, developers may apply it inconsistently, and it increases gas costs. Solidity eliminates the need for SafeMath by performing these checks at the compiler level, though it still introduces some runtime overhead [36, 34]. To prevent precision loss, developers use FullMath library [37]. Another prevention strategy in case of integer over/underflow is using suitable data types [36]. Using formats with sufficient capacity, such as uint256 instead of uint8 helps prevent integer over/underflow issues.
III-A2 Improper Handling of External Calls
Also known as “Mishandled Exception”, “Exception Disorder”, or “Unchecked Call Return Value”, this vulnerability arises when a contract ignores the boolean success flag returned by a low‑level external call. In the EVM the operations call, delegatecall, staticcall, callcode, and the wrappers send and transfer never propagate a revert from the callee; instead they return a tuple (success, data), where success is true on completion and false if the callee runs out of gas, deliberately reverts, or encounters any other error. Failure to test this flag allows the caller to proceed with an invalid state, potentially locking Ether or corrupting storage.
Two particularly common manifestations are documented in the literature [38, 39]. First, when Ether is transferred with send or transfer, only a ‑gas stipend is forwarded. Any recipient whose fallback or receive function requires more gas will revert, but the sender sees only a false return value. If the caller ignores that flag, the Ether remains trapped in the sending contract, sometimes irreversibly. Second, any low‑level call—whether it uses call, delegatecall, staticcall, or callcode—that proceeds without checking success suffers the same silent‑failure risk: the callee’s error is suppressed, control returns to the caller, and subsequent logic executes under false assumptions.
Prevention/Mitigation
To mitigate vulnerabilities related to external calls, developers should verify the return value of such calls. For example, using require(sendResult, "Transfer failed") after send. Additionally, it is recommended to use transfer instead of send, as transfer automatically reverts the entire transaction in case of failure [38]. Since Solidity , the introduction of the try/catch structure has provided a more robust way to handle errors in external calls. Furthermore, adopting well-established standards such as ERC20 and ERC721, which implement secure transaction mechanisms, is advised [40].
III-A3 Reentrancy
The reentrancy vulnerability occurs when functions are maliciously reentered (e.g., through fallback functions) [41]. If the attacker is able to bypass the validity checks in the target contract, they can reenter the callee function maliciously. The vulnerability arises due to two main reasons: first, a contract’s control flow decision relies on some of its state variables that should be updated by the contract itself, but are not, before calling another contract; and second, there is no gas limit when handing over the control flow to another contract. The Decentralized Autonomous Organization (DAO) attack exploited this vulnerability in . As shown in Figure 3, the original contract includes functions for depositing and withdrawing Ether. The primary issue lies in the withdraw function, where the user’s balance is set to zero after Ether has been sent to their address. This sequence allows an attacker to repeatedly call withdraw before the balance update occurs for multiple unauthorized withdrawals. In the attack contract, the attacker leverages the fallback function to invoke withdraw repeatedly before the balance is adjusted, receiving multiple illegal Ether payments.
Mitigation/Prevention
Two well-established techniques, Check-Effects-Interactions (CEI) [42, 43] and reentrancy guard locks, are commonly used to mitigate this issue. In the vulnerable DepositFunds contract (LABEL:lab:valcon), in line 10, the balance is modified after the external call returns, enabling the malicious reentrancy. In CEI, the contract first verifies conditions (Check), then executes external interactions (Interactions), and finally updates the internal state (Effects). In the improved contract, SecureDepositFunds (LABEL:lab:fixcon), which implements the CEI pattern, the user’s balance is reset to zero before transferring the Ether. This ensures that the internal state is updated first (Effects), followed by external interactions (Interactions). By verifying the conditions initially (Check) and ensuring that no funds remain in the account before initiating the transaction, the contract prevents an attacker from making repeated withdrawals even if they attempt to exploit reentrancy [44].
The second approach involves using a lock variable to prevent multiple executions of a sensitive function at the same time. When withdraw is called, the lock is activated to prevent reentry into the function until execution is complete. Once the function finishes, the lock is released, ensuring that no concurrent executions occur. This mechanism effectively neutralizes reentrancy attacks by blocking repeated calls to the function within the same transaction cycle.
III-A4 Dangerous Delegatecall
This vulnerability gained attention after the Parity hack in 2017, which exploited the EVM opcode delegatecall. Unlike a regular call, delegatecall executes the callee’s bytecode while retaining the caller’s storage, msg.sender, and msg.value (Figure 4). Consequently, if the address provided to delegatecall points to untrusted code (code section of contract B in Figure 4), that code runs with full write access to the caller’s state and can arbitrarily modify storage or even trigger self‑destruction, leading to loss of funds or contract functionality [45, 12].
A critical flaw was disclosed in December 2023 after several projects adopted the meta‑transaction standard ERC‑2771 [46]. Many implementations used delegatecall to let the forwarder library determine _msgSender(), but failed to validate the forwarded payload. An attacker could therefore wrap arbitrary calldata inside a trusted forwarded request and trick the callee into believing that the transaction originated from any address of the attacker’s choosing. The technique was observed in the wild, causing losses of ETH and USDC in separate incidents [47, 48].
Mitigation and Prevention
The safest way to reuse code is to rely on stateless Solidity libraries that are linked at compile time with the using for directive; the bytecode of such a library executes in its own context, so any SSTORE or SLOAD it performs affects only the caller’s storage slot that the in‑line assembly refers to, not the library itself [49]. When dynamic dispatch is unavoidable, restrict delegatecall to logic contracts that you have fully audited and whose bytecode hash is hard‑coded or stored behind an owner‑controlled upgrade mechanism. In practice, this means keeping the implementation address in an immutable variable or in a proxy’s implementation slot that can be changed only by a multisig with onlyOwner or AccessControl [50]. Never allow arbitrary user‑supplied addresses to be the target of delegatecall. Upgradable contracts add a second risk surface: if the new implementation introduces or reorders state variables, every storage slot after the change shifts, corrupting data. To avoid this, adhere to an explicit storage‑layout discipline: reserve gaps (uint256[50] private __gap;) at the end of each contract, append new variables only after those gaps, and use advanced Solidity code differencing tools such as SoliDiffy [51] before every upgrade to verify the intention of the modified codes. Standards that encode these patterns, notably EIP‑2535 (Diamond) and the UUPS proxy, embed storage‑slot constants and upgrade guards that make layout violations detectable on‑chain [52, 53]. Leveraging audited reference implementations such as the OpenZeppelin diamond and UUPS modules eliminates common flaws around delegatecall and proxy upgrades [54] (that use delegatecall to upgrade the contract logic).
III-A5 Unbounded Loops in Dynamic Arrays
When a loop in a smart contract runs over all elements of a dynamic array, the gas cost for executing the loop will be directly related to the number of elements. If the array grows over time or contains many elements, the gas cost may exceed the block gas limit. This can lead to the failure of the function’s execution and result in a DoS attack.
Mitigation/Prevention
III-A6 DoS Attack via Owner Account
Many smart contracts have an owner account that controls the contract. If this account is not adequately protected, attackers may exploit it, which could lead to severe consequences, such as permanently locking Ether in the contract [58].
Mitigation/Prevention
Contracts should 1) restrict access to critical functions, 2) use multi-signature approvals for transactions, and 3) avoid reliance on a single account for key operations [59].
III-A7 State‑Revert
A contract is vulnerable to state‑revert abuse when it reveals the result of a state‑changing operation before the change becomes irrevocable, allowing a caller to decide whether to let the transaction stand. The classic setting is a probabilistic reward dApp that pays one of several prizes: an attacker routes the call through an auxiliary contract, records its own balance, invokes the reward function with (bool ok, bytes data) = address(dApp).call(payload), inspects data to see which prize was drawn, and immediately executes require(false) if the prize is unsatisfactory. Because the revert propagates upward, every write inside the dApp, including token transfers, is rolled back, yet the attacker has already learned the outcome. By repeating this cherry‑picking loop until the desirable branch appears, the adversary collects only high‑value rewards and skews the game’s expected value [60].
Prevention/Mitigation
Effective countermeasures postpone disclosure or make it irreversible. Commit‑and‑reveal schemes write a sealed commitment (keccak256(seed, user)) in one transaction and reveal the seed in a later block, so the outcome is fixed before either party knows it. When genuine randomness is needed, verifiable random‑function services (e.g., Chainlink VRF) provide entropy that cannot be observed in time to trigger a revert. Finally, the reward function can burn gas or charge a non‑refundable fee before emitting any tell‑tale data to eliminate the economic incentive to revert [43].
III-A8 Frozen Ether
Frozen Ether (a.k.a. “locked money”) occurs when a contract can accept ETH but, because its withdrawal logic lives in an external library reached via delegatecall, it cannot release the funds once that library is absent. If the library address is wiped with selfdestruct—intentionally or by accident—the forwarder still receives deposits (no code is needed for that), yet every spend attempt reverts, leaving the balance permanently inaccessible [61, 45].
Mitigation/Prevention
Any contract designed to hold ETH should therefore embed an internal withdrawal function, e.g. using call{value: amount}("") or hard‑code an immutable library address that cannot be destroyed; without such a path, deposits may become irreversible [62].
III-A9 Insecure Randomness
This vulnerability affects contracts that require randomness (e.g., gambling and lottery contracts). These contracts often generate pseudo-random numbers using seeds such as block.number, block.timestamp, block.difficulty, or blockhash. Since these seeds can be influenced by blockchain network participants (validators or miners), malicious actors can manipulate these values to alter the outcome in their favor [63, 64].
For example, the VulnerableLottery [65] contract’s dependence on block.timestamp (Listing LABEL:lst:lottery) and a private seed for random number generation results in unsafe randomness. Given that miners can alter block.timestamp within a narrow range, an attacker can affect the result by sending transactions at particular times. Moreover, using a predictable seed does not generate enough entropy to allow for the prediction of future outcomes.
Prevention/Mitigation
For secure random number generation, using blockchain-internal sources like block.timestamp and block.difficulty is insecure due to their transparency and predictability [66]. Instead, developers can utilize external oracles such as Chainlink VRF (Verifiable Random Function), which provides more tamper-resistant randomness [67, 68]. Another approach is the commit-reveal scheme, which enhances randomness security by requiring participants to submit a concealed value first and reveal it later, preventing attackers from predicting or manipulating results [69, 43]. Additionally, employing public-key cryptosystems like the Signidice algorithm [70], originally developed for EOS blockchain and adopted in projects like DAO.Casino, can generate random numbers in two-party contracts. Furthermore, cross-chain oracles such as BTCRelay [71] provide a trustless mechanism for Ethereum smart contracts to verify Bitcoin transactions. In order to enable the use of Bitcoin’s Proof-of-Work-derived entropy as a source of randomness in Ethereum contracts, BTCRelay stores Bitcoin block headers on Ethereum and permits contracts to check Bitcoin transaction inclusion proofs. When combined with other protections against validator manipulation, such as time-delay and multi-source entropy, this strategy is useful.
III-A10 Authorization Using tx.origin
The global tx.origin holds the externally owned account (EOA) that initiated the entire transaction chain, whereas msg.sender refers to the immediate caller of a function. Using tx.origin in access control checks rather than msg.sender lets an attacker insert a malicious contract into the call flow and pass the check as the original EOA.
In LABEL:lst:victim, the Victim.withdrawAll function authorizes withdrawal with require(tx.origin == owner). The attacker contract in LABEL:lst:attacker simply calls withdrawAll; because tx.origin still equals the owner’s EOA, the transfer succeeds and forwards funds to the attacker. No owner approval occurs.
Prevention/Mitigation
Developers should avoid using tx.origin for authorization and instead rely on msg.sender, which represents the immediate caller of the function [72]. Furthermore, it can be beneficial to restrict function access to approved users [73] by putting strong access control mechanisms such as OpenZeppelin Ownable contract.
III-A11 Front‑Running (Abusing Transaction‑Ordering Dependency)
Front-running is a common attack vector that exploits the public and sequential nature of blockchain transaction processing, allowing an adversary to profit by preempting a victim’s pending operation. Because miners (or validators) decide the execution order of pending transactions, an adversary who foresees a profitable state change can insert their own call ahead of the victim’s. This is done either by rebroadcasting the same call with a higher gas price, which miners typically favor [12], or if the adversary controls the block proposer by unilaterally reordering transactions regardless of fees [74]. Such re‑ordering can alter contract outcomes and extract value from honest users. In a simple scenario of front-running attack, Alice broadcasts transaction with gas price at time . Bob observes in the mempool and immediately submits with a higher gas price at . Validators, who generally sort pending transactions by gas price, include first in block ; is confirmed later (in the same or a subsequent block). By pre‑empting the state change encoded in (e.g., a profitable DEX swap), Bob captures its economic value, an archetypal front‑running attack that exploits transaction‑ordering dependency.
Prevention/Mitigation
To address this vulnerability, developers can use techniques like batch auctions, which process several transactions at once to limit manipulation of transaction ordering [75], and commit-reveal schemes, which divide transactions into two phases to initially conceal sensitive information and reduce front-running risk [75].
III-A12 Prodigal Contract
This vulnerability refers to smart contracts that wrongly send Ether to arbitrary addresses and insufficiently control Ether transfers. It is also known in literature as “leaking ether to arbitrary addresses”. During runtime, sending Ether to addresses other than contract’s owners or addresses that have not deposited Ether to the contract, could flag the existence of this vulnerability in a contract [76].
Prevention/Mitigation
Thorough input validation and stringent access rules are necessary to stop prodigal contracts. Unintentional asset leaks are decreased via input validation, which makes sure parameters like recipient addresses or transfer amounts follow certain rules (such as whitelisting addresses or confirming non-zero values) [36]. Another security measure against unauthorized activities is added by restricting important functions (e.g. fund transfers) to authorized entities through the use of Role-Based Access Control (RBAC) or OpenZeppelin’s onlyOwner modifier [54].
III-A13 Ether Lost to Orphan Address
To initiate an Ether transfer, a valid 160-bit address must be specified. If the provided address is invalid or does not exist, the transferred funds will be permanently lost [39]. If the to address is set to the contract’s own address, the transferred tokens will get locked inside the contract with no way to withdraw them, resulting in a loss. This happens because the function does not verify whether the to address is a valid recipient, including ensuring that it’s not the contract itself (see LABEL:lst:transfer).
Prevention/Mitigation
Developers should assure (via require statements) recipient addresses are neither the zero address (address(0)) nor the contract’s own address (address(this)) and confirm the recipient’s capacity to manage transfers (e.g., avoiding non-withdrawable contracts) [54]. Adopting “pull-over-push” withdrawal mechanisms, where users initiate transfers rather than relying on direct pushes, also helps reducing risks of sending to invalid addresses [77].
III-A14 Untrustworthy or Manipulable Data Feeds
Smart contracts often depend on external facts such as token prices, weather data, sports scores to make irreversible decisions. An oracle is the component that transports those facts on chain (or acts as an service contract to provide data from other on-chain services). Recent studies show that integrity of the data can be manipulated during different stages such as creation, forwarding, election, consumption during its lifecycle [4]. We identified two settings where data feeds are manipulatable:
Decentralized Exchange (DEX)‑based on‑chain oracles. System model: a contract queries the current price shown by an Automated Market Maker (AMM) pool such as Uniswap [78]. Attacker model: participants may inject capital (e.g., via a flash loan) to shift the pool’s reserves for a single block and read back a distorted price before arbitrage restores equilibrium [79].
Centralised off‑chain oracles. System model: one or a few trusted reporters sign values off-chain and publish them on-chain. Attacker model: compromise, bribe, or simply delay any reporter to feed stale or false data [80].
Prevention/Mitigation
The literature proposes the following defense strategies:
Robust temporal aggregation. Use multi‑block medians or long‑window time-weighted average pricing (TWAP) taken to smooth out single‑block distortions [4].
Grace periods for finality. A contract should ignore data that has not yet cleared a dispute window to block contested updates.
On‑chain invariants and circuit breakers. Simple require or assert statements that cap per‑block price changes, or tool‑generated invariants such as those from Trace2Inv [81] can stop attacks at runtime.
Based on the mentioned two system settings, combining smoothing, economic incentives, program invariants, and runtime monitoring can raise the adversary’s required capital, coordination, and technical difficulty beyond profitable levels.
III-A15 Compiler Version Not Fixed
This vulnerability occurs when a contract uses an outdated or broadly specified compiler version [85]. For example, specifying pragma solidity ^0.4.0; allows compilation with any version from 0.4.0 to 0.5.0 (excluding 0.5.0). This makes it possible for the contract to be compiled with versions known for certain vulnerabilities. Here, in versions before 0.4.22, if a function intended as a constructor did not match the contract name exactly (e.g., after renaming the contract but not the function), it was treated as a regular public function, not a constructor. This allowed attackers to invoke it maliciously after deployment. By fixing the compiler version explicitly (e.g., pragma solidity 0.4.25;), the contract becomes compilable only with 0.4.25, where the constructor keyword is mandatory, thus mitigating this vulnerability [12].
Prevention/Mitigation
III-A16 Event-Ordering Bug
If distinct sequences of transactions or function invocations lead to an unexpected behavior or final state (post-transactions), it is identified as an EO bug [87].
Prevention/Mitigation
III-A17 Type Casting
This vulnerability stems from improper type conversions in Solidity, such as truncation bugs when casting a larger integer type (e.g., uint16) to a smaller one (e.g., uint8), leading to data loss. Additionally, converting between signed and unsigned types of the same width can cause “signedness bugs,” where negative values turn into large positive ones or vice versa [88, 34].
Prevention/Mitigation
To prevent type-casting vulnerabilities, developers should ensure proper type validation by explicitly checking types and validating value ranges before performing conversions. This reduces the risk of unintended truncation [89]. Additionally, Solidity requires explicit casting to help reduce the risk of implicit conversions, especially between signed and unsigned numbers, which might introduce unpredictable behavior [90]. Refactoring the project to the most recent Solidity version is also beneficial as newer versions support type safety features [91].
III-A18 Ponzi Scheme
In Ponzi contracts, vulnerabilities arise (for the users of the Ponzi services) due to low transparency and deceptive promises of high returns, and legal noncompliance. These schemes commonly rely on a constant influx of new investors to create an illusion of profitability, with returns paid to earlier participants using the funds contributed by newer ones, rather than from actual profits. This scheme is unsustainable and leads to losses for later investors when the inflow of new investments slows or stops [92].
Prevention/Mitigation
User education is important and users must avoid them by spotting characteristics of Ponzi schemes, such as unsustainable returns and a lack of transparency (unavailable, unverified, or obscured contract source code) [93].
III-A19 Storage Collision
Storage Collision occurs when two contracts accessing the same storage space have a different understanding of the storage layout [94, 95, 96, 97]. This vulnerability is often observed when a proxy upgradability design pattern is used [98]. The proxy upgradability design pattern separates a contract’s state from its logic, allowing the system’s functionality to evolve without losing stored data [43]. In this design, a proxy contract holds the persistent state and delegates calls to an external implementation contract using the delegatecall opcode. Initially, the proxy directs calls to a basic implementation. Later, when new features or patches are required, a new implementation contract is deployed. By updating the proxy’s reference to this new contract, the upgraded logic is activated while the original state (the users’ balances, etc.) remains intact [98, 43]. A specific subcategory of this vulnerability is also known as type confusion. This specific subcategory occurs when the runtime context of the logic contract (e.g., in an upgradable contract) confuses the type of an object in storage with another type, leading to its incorrect utilization. For example, when a variable is allocated with one type and later accessed with an incompatible type [94, 99].
For example, the smart contract setup in LABEL:lst:collision is vulnerable to storage collision due to the improper alignment of storage slots between the Proxy and Logic contracts when using delegatecall. Specifically, both contracts store critical state variables in slot 0x0, but with different interpretations: the Proxy contract uses it for visits, while the Logic contract stores initialized and admin in the same slot. When initialize() is called via delegatecall, it writes true (1) to initialized and sets admin to msg.sender, unknowingly overwriting the visits variable in the Proxy contract. An attacker can exploit this by re-invoking initialize(), since the overwritten storage allows it to be called again, setting their own address as admin. Once this is done, the attacker can execute withdraw() to drain the contract’s funds. See LABEL:lst:collision for more details.
Prevention/Mitigation
To avoid storage collisions in proxy patterns, keep the implementation’s storage layout consistent with the proxy’s reserved slots (or assign fixed slots with inline assembly, as in EIP‑1967) [100]. Supplement this with exhaustive tests and fuzzers like Harvey to catch any remaining overlaps pre‑deployment [96].
III-A20 Business Logic Flaws
In smart contracts, business logic flaws (aka accounting errors) occur when financial calculations are inconsistent, inaccurate, or disorganized because of type errors, misinterpreted or mismanaged units (token units), or the implementation is inconsistent with expected behavior by the protocol designer [101].
One such vulnerability happened in Vader protocol [102], where a token unit mismatch occurs when token values are incorrectly combined without the appropriate unit conversion, resulting in an error in the liquidity calculation (calcLiquidityUnits in LABEL:lst:pools). In this case, the scales of the base and token tokens differ. Nevertheless, they are added together in the final formula without the proper conversion, which causes users to lose money and receive an inaccurate liquidity (LABEL:lst:pools).
Prevention/Mitigation
Developers should write extensive test cases covering important business logic scenarios to validate code behavior. This practice helps in identifying discrepancies between the intended and actual functionalities of the smart contract. Extensive testing using model-based testing can deter business logic vulnerabilities [103, 104]. Implementing strict input validation ensures that only correctly formatted and expected data is processed by the contract to prevent unintended behaviors arising from malicious inputs [64].
III-A21 Dangerous Balance Inequality
Input validation and invariant preservation are checked/enforced using assert and require statements in Solidity. Solidity smart contracts on Ethereum alone include more than of these unique checks [105]. In such condition checks, using “==” instead of “” (in statements such as )
when verifying inventory arises from the flawed assumption that the exact match will always occur.
This practice can lead to a vulnerability, as even minor, unexpected changes in inventory values, whether intentional or accidental, could bypass verification logic [85].
Prevention/Mitigation
Developers should use “” to account for potential pre-existing balances [106]. Slither [8], the static analysis tool, provides a “dangerous-strict-equality” detector to identify this vulnerability. Moreover, Continuous Integration (CI) pipelines can automatically check the semantic strength of a pre-/post-condition check in Solidity using SInDi [107].
| Vulnerability | Sec. | Deprecation Reason(s) |
| Upgradable Contracts | §III-B1 | Re-contextualized from a vulnerability to a technical feature. Its risks are manageable through robust on-chain governance and verifiable evolution patterns. |
| Wrong Address | §III-B2 | Mitigated by ABI-decoding requirements in Solidity v0.5.0. Calls with calldata shorter than the expected arguments will revert, preventing the padding that enabled the attack. |
| Erroneous Visibility | §III-B3 | Addressed by Solidity v0.5.0, which mandated explicit function visibility declarations (public, private, etc.). |
| Suicidal Contract | §III-B4 | The behavior of the selfdestruct opcode was severely curtailed by EIP-6780. It now only functions if executed in the same transaction as contract creation, neutralizing its use in common attack patterns. |
| Call-stack Depth Limit | §III-B5 | Rendered economically infeasible by the Tangerine Whistle hard fork (EIP-150) [108], which introduced the “63/64 gas rule.” This rule limits the gas forwarded in a call, preventing deep recursion attacks. |
III-A22 Resource Exhaustion
The gas mechanism in EVM is used to determine how much it costs to execute smart contracts. This system should guarantee that each instruction’s execution cost is calculated according to the resources it uses, including CPU, RAM, and I/O. Nevertheless, some processes, such as blockhash and balance, which require access to blockchain-stored data, are substantially underpriced compared to their actual computational and storage impact due to inconsistent gas pricing for specific EVM instructions [109]. One practical example is looping through a list of accounts to repeatedly retrieve their balance. Since the gas cost for this operation is lower than its actual computational overhead, an attacker can craft transactions that continuously invoke this instruction, causing a significant workload on the network with minimal gas expenditure for the attacker. This shows how inconsistencies in EVM’s resource metering (vs opcode gas) can be leveraged to create transactions that disproportionately consume computational power.
Prevention/Mitigation
Ethereum clients have undergone several upgrades to mitigate resource exhaustion risks. In May 2024, a flaw was exposed in the Geth client’s FeeHistory interface that allowed RPC requests with up to k rewardPercentiles due to missing limits, leading to excessive memory usage and node crashes. The issue was resolved by capping the number of rewardPercentiles [110, 111].
III-A23 Access Control
An access control vulnerability occurs when a smart contract improperly manages access to critical functionalities, allowing unauthorized users to carry out operations including asset transfers, ownership changes, and significant contract settings modifications. Usually, this vulnerability results from the lack of a suitable authentication method. Usual ways to authenticate include implementing Role-Based Access Control (RBAC) [112] using require statements at the entry point of the function body or using a Solidity modifier. This vulnerability is also frequently caused by defining functions as public when only the contract owner or authorized users should be able to access them, failing to implement access control conditions like require(msg.sender == admin) and exposing sensitive variables like owner or balance as public when they should be private to prevent unwanted access [113, 114, 115]. For example, because there is no require(msg.sender == owner, "Not authorized") check in LABEL:lst:accessControl (after line ), any user can alter the contract owner using the public function changeOwner. Likewise, the withdraw function has no access restriction, so anyone can take money out without being authenticated.
Prevention/Mitigation
Developers can utilize OpenZeppelin’s Ownable and AccessControl contracts, which offer role management, to avoid access control vulnerabilities. The onlyOwner modifier [116] in the Ownable contract is used to limit sensitive operations to the owner role. AccessControl in this library enables defining roles and their permissions [117].
III-A24 Timestamp Dependence
This vulnerability occurs when the contract’s logic depends on the value of block.timestamp (the block’s timestamp). This vulnerability arises due to the clock drift of distributed systems, which allows miners to adjust the timestamp within a specific range, typically seconds in Ethereum [118]. Smart contract condition checks may be impacted by this manipulation, even if they have nothing to do with producing random numbers. For instance, consensus participants can alter the end time of events like bids or auctions to their advantage [119]. Validators can reduce or lengthen waiting periods or lock-up periods before specific activities are initiated in a contract, changing the outcome to the advantage of a particular group [43]. Additionally, it is a typical manipulation for miners to reorganize and modify the timing of transactions, particularly in contracts that depend on the precise timing or order of transactions. Changing the timestamp may also alter the outcome of computations that depend on time, causing the contract to behave inconsistently. Miners may modify the voting schedule and outcomes to benefit themselves or a certain group, which could have an effect on governance ecosystems [120].
Prevention/Mitigation
III-B Deprecated Vulnerabilities
| ID | Incident | Date | Loss (USD) | Ecosystem | Application | Vulnerability | Network |
| 1 | Euler Fin. [123, 124, 25] | 2023-03-13 | DeFi | Lending | Access control, Business Logic Flaws | Mainnet | |
| 2 | Nomad Bridge [26, 125] | 2022-08-02 | Bridge | Cross-chain | Access Control | Mainnet | |
| 3 | BonqDAO [126] | 2023-02-03 | DeFi | Stablecoin | Price Manipulation | Sidechains | |
| 4 | WooFi [127, 128] | 2024-03-06 | DeFi | DEX | Price Manipulation | L2 | |
| 5 | Fei Rari [129, 130] | 2022-05-01 | DeFi | Lending | Reentrancy | Mainnet | |
| 6 | Infini [131, 132] | 2025-02-24 | DeFi | Stablecoin | Access Control | Mainnet | |
| 7 | KyberSwap [133, 134] | 2023-11-23 | DeFi | DEX | Business Logic | Mainnet | |
| 8 | Penpie [135, 136] | 2024-09-03 | DeFi | Yield Aggregator | Reentrancy | Mainnet | |
| 9 | Sonne Finance [137, 138] | 2024-05-16 | DeFi | Lending | Rounding Error | L2 | |
| 10 | Inverse Fin.[139, 140, 141] | 2022-04-02 | DeFi | Lending | Price Manipulation | Mainnet | |
| 11 | Holograph [142, 143] | 2024-06-13 | Token | NFT | Access Control | Mainnet | |
| 12 | Deus DAO [144, 145] | 2023-05-05 | DeFi | Token | Access Control | L2 | |
| 13 | Ionic [146, 147] | 2025-02-04 | DeFi | Lending | Input Validation | L2 | |
| 14 | Ronin Bridge [148, 149] | 2024-08-06 | Bridge | Cross-chain Bridge | Access Control | Mainnet | |
| 15 | LI.FI [150, 151, 152] | 2024-07-16 | DeFi | Cross-chain Bridge | Access Control | Mainnet | |
| 16 | Yearn Finance [153, 154] | 2023-04-13 | DeFi | Yield Aggregator | Misconfiguration | Mainnet | |
| 17 | zkLend [155, 156] | 2025-02-12 | DeFi | Lending | Rounding Error | L2 | |
| 18 | LI.FI Protocol[157, 158] | 2024-07-16 | DeFi | Cross-Chain | Unchecked Call Return Value, Access Control | Mainnet | |
| 19 | Hundred Finance [159, 160] | 2023-04-15 | DeFi | Lending | Exchange Rate Manipulation, Rounding Error | L2 | |
| 20 | Abracadabra Money[161] | 2024-01-31 | DeFi | Lending | Business Logic, Rounding Error | Mainnet | |
| 21 | Lodestar Finance [162, 163] | 2022-12-10 | DeFi | Lending | Price Manipulation | L2 | |
| 22 | 1Inch [164, 165] | 2025-03-06 | DeFi | DEX | Integer Underflow | Mainnet | |
| 23 | Shezmu [166, 167] | 2024-09-20 | DeFi | Lending | Access Control | Mainnet | |
| 24 | Gamma Strategies[168, 169] | 2024-01-04 | DeFi | Liquidity Pool | Price Manipulation | L2 | |
| 25 | Conic Finance [170, 171] | 2023-07-25 | DeFi | Lending | Read-Only Reentrancy, Price Manipulation | Mainnet | |
| 26 | KiloEx [172, 173] | 2025-04-17 | DeFi | DEX | Access Control, Price Manipulation | L2 | |
| 27 | SIR Trading [174, 175] | 2025-03-30 | DeFi | Synthetics | Storage Collision | Mainnet | |
| 28 | Abracadabra Money[176] | 2025-03-26 | DeFi | Lending | Business Logic | L2 | |
| 29 | DeltaPrime [177, 178] | 2024-11-11 | DeFi | Lending | Improper Input Validation | L2 | |
| 30 | Onyx Protocol [179, 180] | 2024-09-26 | DeFi | Lending | Exchange Rate Manipulation, Access Control | Mainnet | |
| 31 | Dexible [181, 182] | 2023-02-20 | DeFi | DEX Aggregator | Access Control | Mainnet | |
| 32 | Dough Finance [183, 184] | 2024-07-12 | DeFi | Lending | Improper Input Validation | Mainnet | |
| 33 | CloberDEX [185, 186] | 2024-12-10 | DeFi | DEX | Reentrancy | L2 | |
| 34 | Tornado Cash [187, 188] | 2023-05-20 | DeFi | Privacy Mixer | Governance | Mainnet | |
| 35 | Team Finance [189, 190] | 2022-10-27 | DeFi | Token Locker | Input Validation | Mainnet | |
| 36 | Raft Protocol [191, 192] | 2023-11-10 | DeFi | Stablecoin | Rounding Error | Mainnet | |
| 37 | Rho Market [193, 194] | 2024-07-19 | DeFi | Lending | Priviledge Abuse, Price Manipulation | L2 | |
| 38 | UwuLend [195, 196] | 2024-06-10 | DeFi | Lending | Price Manipulation | Mainnet | |
| 39 | Velocore [197, 198] | 2024-06-02 | DeFi | AMM | Underflow | L2 | |
| 40 | Curio [199, 200] | 2024-03-23 | DeFi | DAO voting system | Access Control, Dangerous Delegatecall | Mainnet | |
| 41 | Unizen [201, 202] | 2024-03-08 | DeFi | DEX | Dangerous Delegatecall | Mainnet | |
| 42 | Seneca Protocol [203, 204] | 2024-02-28 | DeFi | CDP | Input Validation | Mainnet | |
| 43 | OKX DEX [205, 206] | 2023-12-12 | DeFi | DEX | Access Control, Dangerous delegatecall | Mainnet | |
| 44 | Zunami Protocol [207, 208] | 2023-08-13 | DeFi | Stablecoin | Price Manipulation | Mainnet | |
| 45 | EraLend [209, 210] | 2023-07-25 | DeFi | Lending | Read-Only Reentrancy, Price Manipulation | L2 | |
| 46 | Atlantis Loans [211, 212] | 2023-06-10 | DeFi | Lending | Governance | Sidechains | |
| 47 | Sturdy Finance [213, 214] | 2023-06-12 | DeFi | Lending | Read-only Reentrancy, Price Manipulation | Mainnet | |
| 48 | Jimbo’s Protocol [215, 216] | 2023-05-28 | DeFi | Lending/DEX | Access Control, Price Manipulation | L2 | |
| 49 | Swaprum [217, 218] | 2023-05-19 | DeFi | DEX | Access Control | L2 | |
| 50 | Orion Protocol [219, 220] | 2023-02-03 | DeFi | DEX Aggregator | Reentrancy | Mainnet |
This section discusses vulnerabilities in smart contracts that were formerly seen as security issues but have been fixed by protocol enhancements and governance systems. Table IV presents a list of these deprecated Ethereum smart contract vulnerabilities and their brief deprecation reasons.
III-B1 Upgradable Contracts
Smart contracts often incorporate upgradability mechanisms through registries, logic contracts, and proxies to manage immutability constraints of the code section of the contract [221]. Although early analyses classified upgradability as a potential vulnerability [12], recent advancements and governance mechanisms have demonstrated that upgradability is better understood as an inherent technical challenge rather than a security flaw. On-chain governance proposals [222, 223] now enable decentralized control over upgrades, reducing the risk of unilateral malicious actions. Moreover, research efforts such as SoliNomic [224] have introduced formal refinement-based approaches to structuring the evolution of decentralized organizations (for correct and consistent upgrades). Therefore, upgradability, when properly designed, enhances the adaptability of decentralized systems without compromising security.
III-B2 Wrong Address
This vulnerability arises due to the improper handling of function arguments. EVM expects input arguments to be encoded in chunks of 32 bytes; however, if an attacker deliberately sends a truncated address (i.e., an address with missing bytes at the end), EVM automatically pads the missing bytes with zeros, potentially altering the intended values of other arguments. This misalignment can cause an attacker to manipulate transaction amounts or other critical parameters in smart contracts [61, 12]. The attacker could craft a transaction with a shortened address, which, when processed, causes the EVM to pad the missing bytes with zeros, leading to a shift in memory alignment where other function arguments, such as transfer amounts, are misinterpreted (e.g., excessive transfers to the attacker’s address). To prevent this vulnerability, developers should implement strict validation to check that the input data matches the expected size. In other words, the contract should explicitly check msg.data length to ensure proper alignment of the arguments. In modern Solidity versions (0.5.0), with its standard ABI decoding, the short address attack is prevented by design [225]. In these modern versions, if an attacker tries to supply a truncated address (resulting in insufficient calldata), the function call will revert before any parameter decoding or auto-padding occurs.
III-B3 Erroneous Visibility
Erroneous visibility vulnerability emerged from Solidity’s default function visibility behavior prior to version 0.5.0. Functions without explicit visibility modifiers automatically became public, exposing critical operations to unauthorized external calls [226]. Implicit behavior often led to unintentional access to state-modifying, fund management, or administrative functionality. Missing visibility specifiers allowed attackers to execute privileged functions in the Parity multisig wallet incident [227].
By requiring explicit visibility declarations for all functions, Solidity version 0.5.0 released in November 2018, fixed this vulnerability [228]. Contracts without visibility specifiers are now rejected by the compiler, turning a runtime vulnerability into a compile-time error. public, private, internal, or external visibility must be explicitly declared by functions.
III-B4 Suicidal Contract
EVM’s selfdestruct allows the contract to destroy itself and send the remaining Ether in the contract to a target address. Since the Dencun upgrade (EIP‑6780, March 2024) selfdestruct no longer removes code or storage and no longer grants a gas refund, unless it is invoked in the same transaction that deployed the contract [229]. Outside that narrow case, it merely transfers the contract’s Ether balance to the specified address, effectively disabling contract “suicide” and the attacks it enabled.
III-B5 Call‑Stack Depth Limit
The EVM caps a call chain at frames. Before EIP‑150 [108] an attacker could, for about gas, create nested calls and cause every subsequent call from a victim contract to fail (return 0). EIP‑150 raised the sub‑call base cost to gas and, more importantly, restricted each call to forwarding at most of its remaining gas. Because the available gas shrinks geometrically, filling the stack now exceeds the block gas limit, rendering the “stack‑depth attack” economically infeasible. The depth cap nonetheless remains, so contracts should always check the return value (or use try/catch) when interacting with external contracts.
IV How Do Real-World Attacks on Ethereum Smart Contracts Align With Academic Vulnerability Classifications?
To answer RQ2, we conducted a comprehensive analysis of significant smart contract exploits that occurred between and early . The collection, selection, and analysis protocol for these attacks is presented in Section II-B. Table V presents the main attributes of these attacks, including date, volume of financial loss, ecosystem, application type, and the exploited vulnerabilities (identified through manual expert analysis and based on active vulnerabilities in Section III-A). We present more information of each attack and its attributes (victim contact and attacker addresses, etc.) along with a summary of the technical details of each attack in our accompanying repository333https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents/.
Our insights reveal that while real-world successful attacks involve technical weaknesses that could be attributed to classes of vulnerabilities identified by the literature, there are dominant patterns in the incidents that have been unnoticed by the literature. Access control vulnerabilities and price manipulation emerged as the most frequent and financially devastating attack vectors, accounting for a combined loss of $ million in our dataset (see Table VI). Access control-related incidents involve exploit methods beyond mere coding flaws (e.g., lack or misplacement of require(msg.sender==owner) checks) and involve human-centric factors as well (see Insight IV). The high frequency of price manipulation attacks underscores the critical importance of secure, manipulation-resistant oracles (Insight IV). We also observed that while classic vulnerabilities like reentrancy persist, attackers are evolving their techniques, employing sophisticated variants like read-only reentrancy and targeting new EVM features such as transient storage (Insight IV), opening up novel exploit avenues.
Incidents are rarely single-shot failures
Only 19 of the exploits hinged on merely a code defect; the rest were chains of two or more weaknesses (Insight IV) or contained a human factor enabling the incident. For instance, governance mistakes (on a protocol design level) served as the entry point, which attackers then combined with a price-feed or arithmetic flaw to drain funds.
Human factors are involved
Roughly one-third of the breaches trace back to mis-set upgrade parameters, retained developer keys, or copy-paste mistakes-issues that automated analyses seldom flag (Insight IV). Where the literature focuses on the correctness of implementation, our data show that operational discipline and key management are at least as consequential.
Amplifiers and emergent surfaces
Flash-loan liquidity reliably magnified price manipulation attacks (Section IV); protocol composability let a rounding bug in one contract propagate across lending markets (Insight IV); and brand-new EVM features such as transient storage introduced fresh attack vectors (Insight IV). These observations motivate the tiered root-cause model introduced later in Section V.
Insights 1–10 present the identified patterns in detail.
| Vulnerability Category | Number of Incidents | Total Losses (USD) |
| Access Control | 13 | 417,950,000 |
| Price Manipulation | 13 | 279,750,000 |
| Reentrancy | 7 | 118,501,279 |
| Business Logic Flaws | 5 | 177,400,000 |
| Input Validation | 6 | 41,850,000 |
| Rounding Error | 4 | 36,900,000 |
| Integer Underflow | 2 | 11,800,000 |
| Governance | 2 | 4,673,500 |
| Dangerous Delegatecall | 2 | 4,800,000 |
| Storage Collisions | 1 | 355,000 |
For a detailed review report of each incident, refer to our accompanying repository444https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents/ providing in-depth analysis.
V A Multi-Tier Root-Cause Framework for Smart Contract Exploits
We started by identifying smart contract vulnerability classes from academic literature (in Section III) to understanding which classes of these vulnerabilities are indeed exploited in real-world attacks (in Section IV). Our analysis of these high-impact incidents revealed that financially severe incidents often stem from sources beyond simple developer mistakes. These exploits frequently originate from flawed economic designs, broken operational procedures (e.g., governance protocols), and incorrect assumptions about a contract’s environment. Based on the observations from the real-world incident analyses and our comprehensive SLR of vulnerability classifications, we propose a novel, four-tier root-cause-based framework.
The four-tier (not to be confused with blockchain layers) consist of: (1) Flawed Economic Design & Protocol Logic, where the core concept is unsound; (2) Protocol Lifecycle & Governance Failures, where human processes like upgrades, key management, or governance are broken; (3) External Dependency Vulnerabilities, where the protocol makes unsafe assumptions about external data or contracts; and (4) Implementation-Level Weakness, which encompasses classic coding errors. This model reveals that the most significant financial losses often stem from failures in the first two levels, areas that code-centric audits may overlook. This entails that in case of an existing flaw in an early tier (tier 1 and 2), the protocols will be exploitable regardless of quality of coding practices (lack of tiers 3 and 4 attack vectors). This model organizes exploit origins from the most fundamental (the design of the protocol) to the most specific (the implementation). It recognizes that a contract can be technically sound yet economically flawed, or developed securely from known smart contract vulnerabilities but managed insecurely. Table VII summarizes this framework, linking each category to the real-world incidents from our dataset of attacks that exemplify it.
| Tier | Sub-Category | Description Incidents Involved (ID) |
| T1: Flawed Economic Design & Protocol Logic | Incorrect Pricing/Valuation Models | The protocol’s core formulas for calculating value are inherently manipulable. Incidents: 4, 7, 19, 24, 30, 44, 48 |
| State Inconsistency & Desynchronization | The logic allows different parts of the protocol to hold contradictory state information, enabling exploits. Incidents: 20, 28, 45, 47 | |
| Perverse Incentive Mechanisms | The design unintentionally rewards behavior that harms the protocol, such as self-liquidation. Incidents: 1 | |
| T2: Protocol Lifecycle & Governance Failures | Clumsy Deployments & Upgrades | Mistakes during contract initialization or upgrades create vulnerabilities. Incidents: 2, 14, 16, 41 |
| Compromised Keys & Insider Threats | Administrative keys are stolen or abused by internal actors (current or former) to bypass security. Incidents: 6, 11, 37, 43, 49 | |
| Governance System Exploitation | The rules of the DAO are manipulated to pass malicious proposals or seize administrative control. Incidents: 34, 40, 46 | |
| T3: External Dependency Vulnerabilities | Oracle & Price Feed Manipulation | The protocol uncritically trusts an external price feed that is susceptible to manipulation. Incidents: 3, 10, 21, 38 |
| Unvalidated User/Contract Inputs | The protocol fails to sanitize or validate data and external calls originating from untrusted users. Incidents: 13, 15, 18, 29, 31, 32, 35, 42 | |
| T4: Impl.-Level Vulnerabilities | Interaction Pattern Failures | The code violates secure interaction patterns (e.g., CEI), enabling reentrancy. Incidents: 5, 8, 25, 33, 50 |
| Arithmetic Errors | The code is susceptible to over/underflow, or precision loss due to EVM integer math. Incidents: 9, 17, 22, 36, 39 | |
| Flawed Access Control Logic | A flaw in the access control implementation (e.g., reversed parameters) allows it to be bypassed. Incidents: 12, 23, 26 | |
| State & Data Handling Errors | The code incorrectly manages data representation or state, leading to issues like storage collisions. Incidents: 27 |
| Defense | Example | Effect on Incidents (2/14/16/41) |
| Invariant‑checked upgrades | Foundry[233]/Echidna[9] test asserting post‑upgrade state (, correct addresses, guard present). | Blocks deployment; test fails on first run. |
| Staged roll‑outs | Deploy first on canary fork 24h observation mainnet. | Limits loss to canary TVL. |
| Operational gates | Multi‑sig + two‑step pause upgrade un‑pause. | Human review spots nonsense state. |
| Runtime monitors | Circuit‑breaker on abnormal withdrawals / mint rates / external calls. | Halts draining within the first block. |
V-A Tier 1: Flawed Economic Design & Protocol Logic
Here, the protocol is vulnerable not because of a bug, but because its core concept is economically or logically unsound ( incidents trace back to these vulnerabilities). The code may execute exactly as intended, but the intended functionality itself is flawed. For example, the academic vulnerability of business logic flaws (where the contract does not match expected behavior) (Section III-A20) is a subset of this tier. However, our findings show this category is broader than business logic flaws, including the flawed liquidation incentive model in the Euler Finance [24] case and the insecure pricing formula in the WooFi [127] exploit, which are design-level errors, not simple implementation bugs.
V-B Tier 2: Protocol Lifecycle & Governance Failures
Our empirical data reveals that a significant number of catastrophic incidents ( incidents) originate from failures in the human and operational processes surrounding a smart contract. This tier is almost entirely absent from traditional, code-centric vulnerability taxonomies, yet it is a dominant cause of failure. Incidents like the Nomad Bridge exploit, caused by a clumsy upgrade procedure [26], or the Infini exploit, resulting from a failure to revoke developer privileges [132], would not have been detected by existing code analysis tools. They represent failures in deployment, maintenance, and governance procedures that show the need to audit not just the code, but also the team’s operational security.
V-C Tier 3: External Dependency Vulnerabilities
This tier addresses vulnerabilities and incidents arising from a contract’s data dependencies with its environment (12 incidents). The contract’s internal logic may be sound, but it makes unsafe assumptions about the inputs it receives from the outside world. This expands upon the literature’s concept of Untrustworthy Data Feeds (Section III-A14) to include any external dependency. This includes not only price oracles [4], as seen in the Inverse Finance attack [234], but also unvalidated calldata from users, as exploited in the LI.FI incident [150]. Furthermore, our analysis identifies the critical role of environmental factors like the availability of flash loans, which, while not a vulnerability themselves, act as powerful amplifiers for exploits in this tier.
V-D Tier 4: Implementation-Level Weakness
This tier aligns most closely with the traditional vulnerability classifications identified in our SLR (RQ1 results in Section III). We traced back the origin of incidents to these vulnerabilities (Table VII). These are concrete programming mistakes where the contract contains a known vulnerability. This includes well-documented issues such as reentrancy (Section III-A3), as seen in many attacks such as Fei Rari exploit [235], and integer over/underflow (Section III-A1), found in multiple incidents such as Velocore [197]. While critically important, the number of incidents that their root-cause is attributed to other tiers (see Table VII) suggests an incomplete picture of real-world attacks solely based on tier 4 vulnerabilities.
V-E A Case Study for Tier 2
In order to investigate the usefulness of the framework proposed, we analyze “clumsy deployments and upgrades”–a sub-category of tier 2 vulnerabilities (from Table VII). This sub-category consists of contract upgrades and initialization that lead to vulnerabilities. Four of the largest bridge, vault, and DEX catastrophes in our dataset (IDs 2, 14, 16, 41; $M combined loss) share the same root‑cause: a seemingly flawless deployment and upgrade that left the contract in an unsafe state while the byte‑code itself remained bug‑free. Treating them jointly shows that our sub‑category is the minimal sufficient grain at which attacker assumptions, exploit steps, and defenses can be specified once and reused to prevent future exploits to similar protocols.
V-E1 Attacker Model
Based on our incident analyses, we envision an attacker with the following attributes:
Goal: Extract maximum value by exploiting mis‑configured contract state after upgrade or deployment.
Capabilities: i) Craft arbitrary on‑chain transactions (privileged keys not required). ii) Inspect contract storage. iii) Use flash loans to amplify the effect. iv) Reads the verified source code of the protocol/ABI.
Aussumption: Upgrades already live as contracts on mainnet.
V-E2 Exploit Walk‑Through
The incidents can be broken down into three steps:
Step A (State discovery): Attacker reads mis‑initialised variable (Nomad committedRoot, Ronin _totalOperatorWeight, Yearn vault basket, Unizen call‑guard) and confirms it lifts a safety check.
Step B (Trigger): Crafts transaction exploiting the lifted check: replay any root (Nomad), sign with tiny weight (Ronin), mint vault shares (Yearn), or route trade through the attacker contract (Unizen).
Step C (Extraction): Drains bridge funds / swaps inflated shares / spends user approvals. No reentrancy or oracle tricks are needed; the upgrade itself is the point of failure.
V-E3 Defense Model
Table VIII presents our proposed defenses against the attacker model of this sub-category. Before moving real assets, effective tier 2 security must wrap the entire upgrade life‑cycle: (i) pre‑deployment/upgrade initiation property‑based tests that abort if critical state is mis‑initialised; (ii) a canary rollout that risks only a capped pilot TVL while the new code faces live mainnet conditions for h; (iii) human multi‑sig gating that forces a pause‑upgrade‑unpause ritual; and (iv) on‑chain circuit‑breakers [43] that freeze abnormal withdrawals, mint rates, or delegate‑calls within the first block.
VI Conclusion and Future Work
Threats to Validity
Our evidence base contains only incidents that were publicly revealed and financially substantial, so silent failures and minor incidents remained outside the scope of this work which can affect our results. All data came from the EVM blockchain family, which could lead to an alternative execution platform exposing additional incident patterns not observed in our study.
Future Directions
At the economic layer, the community needs agent‑based simulation, model‑checking, and other formal‑methods machinery that can prove the designed system preserves economic soundness. At the governance and lifecycle tier, development pipelines should integrate upgrade simulators, proposal‑linting tools and role‑minimisation checks. External data dependency analysis can evolve into whole‑mempool and cross‑protocol reasoning that surfaces oracle manipulation or flash‑loan amplification before execution. Running across all tiers, real‑time monitoring observing mempool based on exploit‑chain-aware rules should be investigated. Finally, as bridges and roll‑ups are common among loss statistics (Table V), they require more investigation as the contracts interacting are heterogeneous [231].
Conclusion
This study unites the literature’s focus on implementation bugs with the practitioner’s day‑to‑day experience of successful attacks on Ethereum‑based protocols. Through a systematic literature review of peer‑reviewed papers and dissecting fifty publicly disclosed exploits that caused a total loss of $ between 2022 and 2025, we show that less than one‑third of recorded losses can be traced to an isolated coding mistake. The remainder cascaded from flawed incentive design, clumsy governance decisions, unsafe assumptions about external dependencies, or exploit chains-factors that mainstream static and dynamic analyzers rarely take into account. To capture this, we introduced a four‑tier root‑cause framework ranging from economic logic at tier 1 to programming mistakes at tier 4; together with our labeled incident dataset and mitigation catalog, the framework equips auditors, tool builders and protocol designers with a map of the modern threat surface based on real incidents.

### v2

@@
-SoK: Root Cause of $1 Billion Loss in Smart Contract Real-World Attacks via a Systematic Literature Review of Vulnerabilities
+SoK: Root Causes of $1 Billion Loss in Smart Contract Real-World Attacks via a Systematic Literature Review of Vulnerabilities
 Abstract
-The Ethereum ecosystem, despite its maturity, continues to witness catastrophic attacks, with billions of dollars in assets lost annually.
-In response, a significant body of research has focused on identifying and mitigating smart contract vulnerabilities.
-However, these efforts predominantly focus on implementation-level bugs, leaving a critical gap between academic understanding of vulnerabilities and the root causes of real-world high-impact financial losses.
-We employ a two-pronged methodology: first, a systematic literature review of 71 academic papers to build a comprehensive and up-to-date catalog of 24 active and 5 deprecated vulnerabilities as understood by the research community.
-Second, we conduct an in-depth, empirical analysis of 50 of the most severe real-world exploits between 2022 and 2025, collectively incurring over $1.09B in losses, to identify their true root causes.
+While catastrophic attacks on Ethereum persist, vulnerability research remains fixated on implementation-level smart contract bugs, creating a gap between academic understanding of vulnerabilities and the true root causes of high-impact, real-world incidents.
+To address this, we employ a two-pronged methodology: first, a systematic literature review of 71 academic papers to build a catalog of 24 active and 5 deprecated vulnerabilities.
+Second, we conduct an in-depth, empirical analysis of 50 of the most severe real-world attacks between 2022 and 2025, collectively incurring over $1.09B in losses, to identify their root causes.
 We introduce the concept of “exploit chains” by revealing that many incidents are not caused by isolated vulnerabilities but by combinations of human, operational, and economic design flaws that link with implementation bugs to enable an attack.
-Our analysis yields insights on how DApps are exploited in practice, leading to a novel, four-tier root-cause framework that moves beyond code-level vulnerabilities.
-We find that real-world successful attacks on Ethereum (and related networks) trace back to one of the four tiers of (1) protocol logic design, (2) lifecycle and governance, (3) external dependencies, and (4) traditional implementation bugs (classic smart contract vulnerabilities).
+Our analysis yields insights on how decentralized applications are exploited in practice, leading to a novel, four-tier root-cause framework that moves beyond code-level vulnerabilities.
+We find that real-world successful attacks on Ethereum (and related networks) trace back to one of the four tiers of (1) protocol logic design, (2) lifecycle and governance, (3) external dependencies, and (4) classic smart contract vulnerabilities.
 We investigate the suitability of this multi-tier incident root-cause framework via a case study.
@@
 However, our review unveils a few gaps in this literature: First, prevention gap: only 4 of 13 works (S4, S9, S12, and S13) address prevention methods, most of which focus exclusively on detection.
-Second, the validation gap: merely 3 works validate their findings against real-world exploits—S3, S9 and S12 or analyze case studies, and exploits, while others rely on synthetic datasets.
+Second, the validation gap: merely 3 works validate their findings against real-world exploits—S3 (case studies), S9 (181 exploits) and S12 (92 exploits), while others rely on synthetic datasets.
 Third, the taxonomy evolution shows a shift from simple categorization to sophisticated multi-layer frameworks (4L, 5L), yet none adopt a root-cause approach based on real incidents.
-Real-World Incidents Tell-Tale
+Telltale Signs from Real-World Incidents
 At 08:56 UTC on 13 March 2023, the Euler Finance lending market was solvent; four minutes later, it was $ million in deficit [24, 25].
@@
 It highlights a systemic issue: while literature focuses on finding bugs in the code, a large-scale analysis of real-world attacks and comparing the exploited vulnerabilities in these incidents and their alignment with the literature is required.
+Thus, we propose the following research questions:
 1.
@@
 2.
-RQ2: What factors, other than mere coding vulnerability patterns (identified in RQ1), are evident in the incidents and contribute to the success of the attack?
+RQ2: What factors, other than mere coding vulnerability patterns (identified in RQ1), are evident in the real-world incidents and contribute to the success of the attacks?
 3.
@@
 II-A Academic Literature Study Protocol
-To answer RQ1, we conducted a systematic literature review (SLR) to obtain a comprehensive and structured understanding of Ethereum smart contract vulnerabilities identified in the academic literature.
-Following Kitchenham’s guidelines [27], our review process includes study identification (Section II-A1) and study selection (Section II-A2).
+To answer RQ1, we conducted a systematic literature review (SLR) to obtain a comprehensive and structured understanding of Ethereum smart contract vulnerabilities from academic literature.
+Following Kitchenham‘’s guidelines [27], our review process includes study identification (Section II-A1) and study selection (Section II-A2).
 II-A1 Study Identification
@@
 Full-text assessment based on availability (EC4) further refined the selection to papers.
-Finally, quality-based evaluation via manual expert empirical validation and identification of novel contributions resulted in papers for in-depth analysis.
+Finally, quality-based evaluation via manual review of the paper and identification of novel contributions resulted in papers for in-depth analysis.
 Figure 1 illustrates this process.
@@
 Mitigation/Prevention
-Contracts should 1) restrict access to critical functions, 2) use multi-signature approvals for transactions, and 3) avoid reliance on a single account for key operations [59].
+Contracts should restrict access to critical functions multi-signature approvals for transactions to avoid reliance on a single account for key operations [59].
 III-A7 State‑Revert
@@
 III-A14 Untrustworthy or Manipulable Data Feeds
-Smart contracts often depend on external facts such as token prices, weather data, sports scores to make irreversible decisions.
+Smart contracts often depend on external facts such as token prices, weather data, and sports scores to make irreversible decisions.
 An oracle is the component that transports those facts on chain (or acts as an service contract to provide data from other on-chain services).
-Recent studies show that integrity of the data can be manipulated during different stages such as creation, forwarding, election, consumption during its lifecycle [4].
-We identified two settings where data feeds are manipulatable:
+Recent studies show that integrity of the data can be manipulated at stages of its lifecycle [4].
+We identified two settings where data feeds of these oracles are manipulable:
 Decentralized Exchange (DEX)‑based on‑chain oracles.
@@
 Attacker model: participants may inject capital (e.g., via a flash loan) to shift the pool’s reserves for a single block and read back a distorted price before arbitrage restores equilibrium [79].
-Centralised off‑chain oracles.
-System model: one or a few trusted reporters sign values off-chain and publish them on-chain.
-Attacker model: compromise, bribe, or simply delay any reporter to feed stale or false data [80].
+Off‑chain oracles.
+System model: trusted reporters sign values off-chain and the oracle publishes them on-chain.
+Attacker model: compromise, bribe, or simply delay reporters to feed stale or false data [80].
 Prevention/Mitigation
 The literature proposes the following defense strategies:
-Robust temporal aggregation. Use multi‑block medians or long‑window time-weighted average pricing (TWAP) taken to smooth out single‑block distortions [4].
+Temporal aggregation. To deter price manipulation in DEX-based oracles, use multi‑block medians or long‑window time-weighted average pricing (TWAP) taken to smooth out single‑block distortions [4].
 Grace periods for finality. A contract should ignore data that has not yet cleared a dispute window to block contested updates.
-On‑chain invariants and circuit breakers.
-Simple require or assert statements that cap per‑block price changes, or tool‑generated invariants such as those from Trace2Inv [81] can stop attacks at runtime.
+On‑chain invariants and circuit breakers. Use invariants (in form of require statements) that cap per‑block price changes to stop attacks at runtime [81] .
 Based on the mentioned two system settings, combining smoothing, economic incentives, program invariants, and runtime monitoring can raise the adversary’s required capital, coordination, and technical difficulty beyond profitable levels.
@@
 III-A24 Timestamp Dependence
-This vulnerability occurs when the contract’s logic depends on the value of block.timestamp (the block’s timestamp).
-This vulnerability arises due to the clock drift of distributed systems, which allows miners to adjust the timestamp within a specific range, typically seconds in Ethereum [118].
+This vulnerability occurs when the contract’s logic depends on the value of block.timestamp.
+Due to the clock drift of distributed systems, miners are allowed to adjust the timestamp within a specific range, typically seconds in Ethereum [118].
 Smart contract condition checks may be impacted by this manipulation, even if they have nothing to do with producing random numbers.
-For instance, consensus participants can alter the end time of events like bids or auctions to their advantage [119].
-Validators can reduce or lengthen waiting periods or lock-up periods before specific activities are initiated in a contract, changing the outcome to the advantage of a particular group [43].
+For instance, consensus participants can alter the end time of events like bids or auctions [119].
+Validators can lengthen waiting or lock-up periods before specific activities to change the outcome to the advantage of a particular group [43].
 Additionally, it is a typical manipulation for miners to reorganize and modify the timing of transactions, particularly in contracts that depend on the precise timing or order of transactions.
@@
 III-B1 Upgradable Contracts
-Smart contracts often incorporate upgradability mechanisms through registries, logic contracts, and proxies to manage immutability constraints of the code section of the contract [221].
-Although early analyses classified upgradability as a potential vulnerability [12], recent advancements and governance mechanisms have demonstrated that upgradability is better understood as an inherent technical challenge rather than a security flaw.
-On-chain governance proposals [222, 223] now enable decentralized control over upgrades, reducing the risk of unilateral malicious actions.
+Smart contracts often incorporate upgradability mechanisms through logic contracts and proxies to manage immutability constraints of the code section of the contract [221].
+Although early analyses classified upgradability as a potential vulnerability [12], recen advancements in upgradability patterns and governance mechanisms have demonstrated that this is better understood as a technical challenge rather than a security weakness.
+On-chain governance proposals [222] now enable decentralized control over upgrades, reducing the risk of unilateral malicious actions.
 Moreover, research efforts such as SoliNomic [223] have introduced formal refinement-based approaches to structuring the evolution of decentralized organizations (for correct and consistent upgrades).
-Therefore, upgradability, when properly designed, enhances the adaptability of decentralized systems without compromising security.
+Therefore, upgradability, when properly designed, enhances the adaptability of DApps without compromising security.
 III-B2 Wrong Address
@@
 Table V presents the main attributes of these attacks, including date, volume of financial loss, ecosystem, application type, and the exploited vulnerabilities (identified through manual expert analysis and based on active vulnerabilities in Section III-A).
-We present more information of each attack and its attributes (victim contact and attacker addresses, etc.) along with a summary of the technical details of each attack in our accompanying repository333https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents/.
-Our insights reveal that while real-world successful attacks involve technical weaknesses that could be attributed to classes of vulnerabilities identified by the literature, there are dominant patterns in the incidents that have been unnoticed by the literature.
+We present more information of each attack and its attributes (victim contact and attacker addresses, etc.) along with a summary of the technical details of each attack in our accompanying repository.
+Our insights reveal that while real-world successful attacks involve technical weaknesses attributed to classes of known vulnerabilities (Section III-A), there are dominant patterns in the incidents that have been unnoticed by the literature.
 Access control vulnerabilities and price manipulation emerged as the most frequent and financially devastating attack vectors, accounting for a combined loss of $ million in our dataset (see Table VI).
-Access control-related incidents involve exploit methods beyond mere coding flaws (e.g., lack or misplacement of require(msg.sender==owner) checks) and involve human-centric factors as well (see Insight IV).
+Access control-related incidents involve exploit methods beyond mere coding flaws (e.g., lack or misplacement of require(msg.sender==owner) checks) and include human-centric factors as well (see Insight IV).
 The high frequency of price manipulation attacks underscores the critical importance of secure, manipulation-resistant oracles (Insight IV).
-We also observed that while classic vulnerabilities like reentrancy persist, attackers are evolving their techniques, employing sophisticated variants like read-only reentrancy and targeting new EVM features such as transient storage (Insight IV), opening up novel exploit avenues.
+We also observed that while classic vulnerabilities like reentrancy persist, attackers are employing more sophisticated variants like read-only reentrancy and targeting new EVM features such as transient storage (Insight IV), opening up novel exploit avenues.
 Incidents are rarely single-shot failures
@@
 | Storage Collisions | 1 | 355,000 |
-For a detailed review report of each incident, refer to our accompanying repository444https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents/ providing in-depth analysis.
+For a detailed review of each incident, refer to our accompanying repository333https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents providing in-depth analysis.
 V A Multi-Tier Root-Cause Framework for Smart Contract Exploits
@@
 | Staged roll‑outs | Deploy first on canary fork 24h observation mainnet. | Limits loss to canary TVL. |
-| Operational gates | Multi‑sig + two‑step pause upgrade un‑pause. | Human review spots nonsense state. |
+| Operational gates | Multi‑sig + two‑step pause upgrade un‑pause. | Human review spots violating states. |
 | Runtime monitors | Circuit‑breaker on abnormal withdrawals / mint rates / external calls. | Halts draining within the first block. |
@@
 This study unites the literature’s focus on implementation bugs with the practitioner’s day‑to‑day experience of successful attacks on Ethereum‑based protocols.
-Through a systematic literature review of peer‑reviewed papers and dissecting fifty publicly disclosed exploits that caused a total loss of $ between 2022 and 2025, we show that less than one‑third of recorded losses can be traced to an isolated coding mistake.
+Through a systematic literature review of peer‑reviewed papers and dissecting publicly disclosed exploits that caused a total loss of $ between 2022 and 2025, we show that less than one‑third of recorded losses can be traced to an isolated coding mistake.
 The remainder cascaded from flawed incentive design, clumsy governance decisions, unsafe assumptions about external dependencies, or exploit chains-factors that mainstream static and dynamic analyzers rarely take into account.